# Rotating metrics and new multipole moments from scattering amplitudes in arbitrary dimensions

### v1

@@ I Introduction
Remarkably, loop corrections in this expansion give rise not only to Planck-suppressed quantum terms, but also to entirely classical terms that survive in the $\hbar\to 0$ limit Iwasaki (1971); Donoghue (1994b); Bjerrum-Bohr et al. (2003); Donoghue and Torma (1996); Holstein and Donoghue (2004); Bjerrum-Bohr et al. (2018); Kosower et al. (2019); Cheung et al. (2018). Specifically, the Schwarzschild metric at first order in a post-Minkowskian (i.e., in powers of the gravitational coupling $G$) expansion can be obtained from the scattering amplitude of a massive scalar field emitting gravitons at tree level, while next order post-Minkowskian corrections are obtained by the classical contributions of each $n-$loop amplitude containing graviton vertices [^1] Recently, the authors of Damgaard and Lee (2024) managed to resum the post-Minkowskian series and obtain the exact Schwarzschild solution. Bjerrum-Bohr et al. (2003). This result generalizes to the case of charged and/or spinning geometries by computing the scattering of particles with electric charge and/or spin, reproducing the post-Minkowskian expansion of the Reissner-Nordstrom, Kerr, and Kerr-Newman metrics in $D=4$ spacetime dimensions Bjerrum-Bohr et al. (2003); Donoghue et al. (2002); D’Onofrio et al. (2022); Gambino (2022).
@@
In this paper we extend this program to the case of spinning geometries both in $D=4$ and $D>4$ dimensions. The motivation for such computation is twofold. First, in $D=4$, unlike the case of spherically symmetric spacetimes, the absence of Birkhoff’s theorem in axisymmetry implies that the vacuum region outside a spinning object is not necessarily described by the Kerr geometry. Within a post-Minkowskian expansion, the leading-order (linear) angular momentum term is universal, but different spinning objects may have different mass quadrupole moments Hartle (1967); Hartle and Thorne (1968); Raposo et al. (2019). Since the spin-induced quadrupole moment is quadratic in the black-hole spin [^2] Note that, in this paper, we occasionally refer to “spin” both for the spin $s$ of quantum fields and for the angular momentum $J$ of the compact object (e.g., a black hole)., in order to extend the analysis of Bjerrum-Bohr et al. (2003) it is necessary to perform computations up to quadratic order in the angular momentum. We shall show that the computation based on the scattering amplitudes provides the post-Minkowskian expansion of the vacuum Hartle-Thorne solution Hartle (1967); Hartle and Thorne (1968), which describes the spacetime of a generic spinning object up to quadratic order in the spin. The Kerr black hole is a particular case of this family, wherein the mass quadrupole moment is fixed by regularity at the horizon and by the black-hole uniqueness theorem Israel (1967); Carter (1971); Hawking (1972); Robinson (1975, 2009); Hawking and Ellis (2011); Heusler (1998); Chrusciel et al. (2012). As we shall discuss, at the level of the effective quantum field theory, the Kerr black hole corresponds to the case of minimal couplings, while any other choice of the quadrupole moment requires non-minimal couplings in the action. The authors of Chung et al. (2019) proved that in $D=4$ the simplest massive $S$-matrix, defined as the term in the 3-point amplitude which behaves well in the UV, reproduces the Kerr metric in limit in which the spin $s\rightarrow+\infty$. Furthermore, the authors of Skvortsov and Tsulaia (2023) proved that a spin-s field minimally coupled to gravity also reproduces the dynamics of Kerr black holes, suggesting a correspondence between the simplest massive $S$-matrix and the minimally coupled action in a QFT description for generic spin $s$. In this sense, one can interpret the minimality of the quantum field theory as the scattering-amplitude counterpart of the celebrated black-hole no-hair theorems in $D=4$ General Relativity Hawking (1972); Robinson (2009); Hawking and Ellis (2011); Chrusciel et al. (2012).
@@
In order to obtain the above results we need to compute the metric from the scattering amplitudes of a massive spin-1 field up to the quadrupole order in a post-Minkowskian expansion. We performed our computations up to two loops (i.e., up to third post-Minkowskian order). The full result is provided in an ancillary Mathematica notebook.
@@ III.2 Non-minimal vertex
In particular, in $d=3$ a tree-level insertion of $\hat{\tau}_{\Phi^{2}h,\text{HL}}^{\mu\nu}$ contributes to 3PM, while in $d=4$ it contributes starting from 2PM.
@@ III.3 Loop amplitudes
From the point of view of the Einstein equations, the divergences discussed above are interpreted as a consequence of the fact that, for a specific gauge choice, imposing that the metric is a series expansion in powers of $\rho(r)$ is not consistent. Similarly, the presence of the free parameters $\Omega_{i}^{\mathrm{free}}$ is interpreted as a redundancy of the gauge defined in Eq. (6). In the next section we will discuss in detail the properties of the metric and give its full expression at 1PM. Although we have also managed to perform such calculation up to 2-loops and for an arbitrary value of $\alpha$, higher orders in the PM expansion do not give any more physical information than the leading order, and since the full expressions are quite involved, we only provide the analytic form of $h_{\mu\nu}^{(2)}(r)$ and $h_{\mu\nu}^{(3)}(r)$ in an auxiliary Mathematica file.
@@ IV.1 Eliminating redundant parameters
In summary, besides the spacetime mass $m$ and spin tensor $S^{\mu\nu}$, the solution depends on two further physical parameters, namely $H_{1}$ and $H_{2}$. However, in $d=3$ this is not the end of the story. Indeed, such case is special since we can write $S^{ij}=\epsilon^{ijk}S_{k}$, where $S^{k}$ is the angular momentum vector and $\epsilon^{ijk}$ is the Levi-Civita symbol. Replacing this relation inside the metric in $d=3$ and fixing
@@ IV.2 Multipole moments
$\mathbb{G}^{(\ell)}_{ij,A_{\ell}}=\tilde{\mathbb{G}}^{(\ell)}_{ij,A_{\ell}}+\frac{1}{2}\delta_{ij}\left(\mathbb{M}^{(\ell)}_{A_{\ell}}-\tilde{\mathbb{G}}^{(\ell)}_{kk,A_{\ell}}\right)\ ,$ (80)
@@ V.1 Hartle-Thorne metric in $D=4$
From the amplitude perspective, this condition is satisfied by an infinite number of non-minimally coupled theories, since $H_{1}$ and $H_{2}$ can be chosen freely as long as the Kerr condition in Eq. (88) is satisfied. This can be associated to the degeneracy occurring in $D=4$, that reduces the number of physical degrees of freedom to one, still having two independent free parameters. In addition to non-minimal QFTs, the minimally-coupled theory, that gives rise to what we define the simplest metric [^12] This name is inspired by the fact that a minimally-coupled theory is the simplest possible theory from a QFT perspective., also satisfies the Kerr condition. Thus, from the point of view of the effective QFT, the scattering amplitude computation unveils that the simplest vacuum solution in $D=4$ is the Kerr black hole.
@@ VI Conclusions
However, our investigation reveals that the situation is different and richer in higher ($D>4$) dimensions. Here, through our scattering amplitude analysis up to 2-loop calculations, we have obtained the generic solution including quadrupole-moments terms quadratic in the object’s angular momentum. Interestingly, at variance with the $D=4$ case, we have found that the solution depends on two quadrupole moment parameters, namely the standard mass quadrupole and a new stress quadrupole; the latter does not exist in $D=4$ and was previously missed in the analysis of Heynen and Mayerson (2023). We have successfully identified and characterized different $D>4$ black-hole solutions, such as the Myers-Perry and the black ring metrics, as particular cases obtained for specific choices of the parameters of our general PM solution. Notably, unlike the four-dimensional case, none of these solutions corresponds to the choice of minimal couplings in the effective action, highlighting that the nature of black holes in higher dimensions is more intricate.
@@ Appendix B Hartle-Thorne metric in harmonic coordinates
Once the coordinate transformation is fixed, we can find the metric in harmonic gauge at the chosen perturbative order by means of (112). We notice that the coordinate transformation has some gauge redundancies since there are coefficients which are unfixed. At 1PM and 2PM the redundancy is parameterized by two coefficients, namely $\mathcal{C}^{(R)}_{1,2,0}$ and $\mathcal{C}^{(R)}_{1,2,2}$. At 3PM a new redundancy arises by means of the unfixed coefficients $\mathcal{C}^{(R)}_{3,0,0}$ and $\mathcal{C}^{(R)}_{3,2,2}$. While the full 3PM result is given in the attached Mathematica notebook, at 1PM the explicit expression of the Hartle-Thorne metric in harmonic gauge reads
@@ Appendix C Myers-Perry black holes in $D=5$ in harmonic coordinates
As we did in the Hartle-Thorne case, replacing Eq. (123) in the harmonic equation, one is able to fix the coefficients of the coordinate transformation. In this case the equivalent of Eq. (115) for the Myers-Perry solution is way more involved, and since there is no physical information in it, we provide it in the attached Mathematica notebook. However, it is worth mentioning that coherently to what we know from the amplitude calculation, the coordinate transformation develops two independent gauge redundancies at orders $O(Ga^{2})$, and two more at orders $O(G^{2}a^{0})$ and $O(G^{2}a^{2})$. In particular the redundancies at 2PM are related to the logarithmic terms and are the classical counterpart of the higher-loop vertex in Eq. (51).
Finally, we report the 3PM Myers-Perry metric in harmonic coordinates in the attached Mathematica file, giving here for the sake of clarity only the expression at 1PM in the case in which $\mathcal{C}_{1,2,0}^{(R)}=0$ and $\mathcal{C}_{1,2,2}^{(R)}=-2/9$ such that $g_{R\Theta}=0$, resulting in the simplified expression
@@ Appendix D Black rings with one angular momentum in harmonic coordinates
Finally, considering the explicit expression of Eq. (130) (reported in the Mathematica file), the black ring metric in harmonic coordinates (in a suitable gauge), in which we are neglecting quadrupoles that are not spin-induced, reads

### v3

@@ I Introduction
Remarkably, loop corrections in this expansion give rise not only to Planck-suppressed quantum terms, but also to entirely classical terms that survive in the $\hbar\to 0$ limit Iwasaki (1971); Donoghue (1994b); Bjerrum-Bohr et al. (2003); Donoghue and Torma (1996); Holstein and Donoghue (2004); Bjerrum-Bohr et al. (2018); Kosower et al. (2019); Cheung et al. (2018); Guevara (2019). Specifically, the Schwarzschild metric at first order in a post-Minkowskian (i.e., in powers of the gravitational coupling $G$) expansion can be obtained from the scattering amplitude of a massive scalar field emitting gravitons at tree level, while next order post-Minkowskian corrections are obtained by the classical contributions of each $n-$loop amplitude containing graviton vertices [^1] Recently, the authors of Damgaard and Lee (2024) managed to resum the post-Minkowskian series and obtain the exact Schwarzschild solution. Bjerrum-Bohr et al. (2003). This result generalizes to the case of charged and/or spinning geometries by computing the scattering of particles with electric charge and/or spin, reproducing the post-Minkowskian expansion of the Reissner-Nordstrom, Kerr, and Kerr-Newman metrics in $D=4$ spacetime dimensions Bjerrum-Bohr et al. (2003); Donoghue et al. (2002); D’Onofrio et al. (2022); Gambino (2022).
@@
In this paper we extend this program to the case of spinning geometries both in $D=4$ and $D>4$ dimensions. The motivation for such computation is twofold. First, in $D=4$, unlike the case of spherically symmetric spacetimes, the absence of Birkhoff’s theorem in axisymmetry implies that the vacuum region outside a spinning object is not necessarily described by the Kerr geometry. Within a post-Minkowskian expansion, the leading-order (linear) angular momentum term is universal, but different spinning objects may have different mass quadrupole moments Hartle (1967); Hartle and Thorne (1968); Raposo et al. (2019). Since the spin-induced quadrupole moment is quadratic in the black-hole spin [^2] Note that, in this paper, we occasionally refer to “spin” both for the spin $s$ of quantum fields and for the angular momentum $J$ of the compact object (e.g., a black hole)., in order to extend the analysis of Bjerrum-Bohr et al. (2003) it is necessary to perform computations up to quadratic order in the angular momentum. We shall show that the computation based on the scattering amplitudes provides the post-Minkowskian expansion of the vacuum Hartle-Thorne solution Hartle (1967); Hartle and Thorne (1968), which describes the spacetime of a generic spinning object up to quadratic order in the spin. The Kerr black hole is a particular case of this family, wherein the mass quadrupole moment is fixed by regularity at the horizon and by the black-hole uniqueness theorem Israel (1967); Carter (1971); Hawking (1972); Robinson (1975, 2009); Hawking and Ellis (2011); Heusler (1998); Chrusciel et al. (2012). As we shall discuss, at the level of the effective quantum field theory, the Kerr black hole corresponds to the case of minimal couplings, while any other choice of the quadrupole moment requires non-minimal couplings in the action. The authors of Chung et al. (2019); Guevara et al. (2019) proved that in $D=4$ the simplest massive $S$-matrix, defined as the term in the 3-point amplitude which behaves well in the UV, reproduces the Kerr metric in limit in which the spin $s\rightarrow+\infty$. Furthermore, the authors of Skvortsov and Tsulaia (2023) proved that a spin-s field minimally coupled to gravity also reproduces the dynamics of Kerr black holes, suggesting a correspondence between the simplest massive $S$-matrix and the minimally coupled action in a QFT description for generic spin $s$. In this sense, one can interpret the minimality of the quantum field theory as the scattering-amplitude counterpart of the celebrated black-hole no-hair theorems in $D=4$ General Relativity Hawking (1972); Robinson (2009); Hawking and Ellis (2011); Chrusciel et al. (2012).
@@
In order to obtain the above results we need to compute the metric from the scattering amplitudes of a massive spin-1 field up to the quadrupole order in a post-Minkowskian expansion. We performed our computations up to two loops (i.e., up to third post-Minkowskian order). The full result is provided in an ancillary Mathematica notebook Anc .
@@ III.2 Non-minimal vertex
In particular, in $d=3$ a tree-level insertion of $\hat{\tau}_{\Phi^{2}h,\text{HL}}^{\mu\nu}$ contributes to 3PM, while in $d=4$ it contributes starting from 2PM. Notice that the $\Omega_{1}$ piece in Eq. (51) exactly corresponds to the counter-term defined in Mougiakakos and Vanhove (2021).
@@ III.3 Loop amplitudes
From the point of view of the Einstein equations, the divergences discussed above are interpreted as a consequence of the fact that, for a specific gauge choice, imposing that the metric is a series expansion in powers of $\rho(r)$ is not consistent. Similarly, the presence of the free parameters $\Omega_{i}^{\mathrm{free}}$ is interpreted as a redundancy of the gauge defined in Eq. (6). In the next section we will discuss in detail the properties of the metric and give its full expression at 1PM. Although we have also managed to perform such calculation up to 2-loops and for an arbitrary value of $\alpha$, higher orders in the PM expansion do not give any more physical information than the leading order, and since the full expressions are quite involved, we only provide the analytic form of $h_{\mu\nu}^{(2)}(r)$ and $h_{\mu\nu}^{(3)}(r)$ in an auxiliary Mathematica file Anc .
@@ IV.1 Eliminating redundant parameters
In summary, besides the spacetime mass $m$ and spin tensor $S^{\mu\nu}$, the solution depends on two further physical parameters, namely $H_{1}$ and $H_{2}$, while every other free coefficient parameterizes a gauge redundancy. This means that in the non-minimal action in Eq. (47) every term can be reabsorbed by a coordinate transformation except for the terms proportional to $K_{3}$ and $K_{4}$, which are physical and are not associated with a gauge transformation. However, in $d=3$ this is not the end of the story. Indeed, such case is special since we can write $S^{ij}=\epsilon^{ijk}S_{k}$, where $S^{k}$ is the angular momentum vector and $\epsilon^{ijk}$ is the Levi-Civita symbol. Replacing this relation inside the metric in $d=3$ and fixing
@@ IV.2 Multipole moments
$\mathbb{G}_{ij,A_{\ell}}^{(\ell)}=\tilde{\mathbb{G}}_{ij,A_{\ell}}^{(\ell)}+\frac{1}{2}\delta_{ij}\Big{(}\mathbb{M}^{(\ell)}_{A_{\ell}}-\tilde{\mathbb{G}}_{kk,A_{\ell}}^{(\ell)}\Big{)}\ ,$ (80)
@@ V.1 Hartle-Thorne metric in $D=4$
From the amplitude perspective, this condition is satisfied by an infinite number of non-minimally coupled theories, since $H_{1}$ and $H_{2}$ can be chosen freely as long as the Kerr condition in Eq. (88) is satisfied. This can be associated to the degeneracy occurring in $D=4$, that reduces the number of physical degrees of freedom to one, still having two independent free parameters. In addition to non-minimal QFTs, the minimally-coupled theory, that gives rise to what we define the simplest metric [^12] This name is inspired by the fact that a minimally-coupled theory is the simplest possible theory from a QFT perspective., also satisfies the Kerr condition. Thus, from the point of view of the effective QFT, the scattering amplitude computation unveils that the simplest vacuum solution in $D=4$ is the Kerr black hole. This is analogous to what was shown in Mougiakakos and Vanhove (2021); D’Onofrio et al. (2022), namely that there exists a reference frame in which the Schwarzschild-Tangherlini metric is induced by the minimally-coupled scalar action.
@@ VI Conclusions
However, our investigation reveals that the situation is different and richer in higher ($D>4$) dimensions. Here, through our scattering amplitude analysis up to 2-loop calculations, we have obtained the generic solution including quadrupole-moments terms quadratic in the object’s angular momentum. Interestingly, at variance with the $D=4$ case, we have found that the solution depends on two quadrupole moment parameters, namely the standard mass quadrupole and a new stress quadrupole; the latter does not exist in $D=4$ and was previously missed in the analysis of Heynen and Mayerson (2023). We have successfully identified and characterized different $D>4$ black-hole solutions, such as the Myers-Perry and the black ring metrics, as particular cases obtained for specific choices of the parameters of our general PM solution. Notably, unlike the four-dimensional case, none of these solutions corresponds to the choice of minimal couplings in the effective action, highlighting that the nature of black holes in higher dimensions is more intricate. In this direction our work could be important to extend the analysis of 4 and 5-point classical gravitational scattering amplitudes involving spinning bodies Siemonsen and Vines (2020); Vines (2018); Vines et al. (2019); Guevara et al. (2019); Luna et al. (2023); Brandhuber et al. (2024); De Angelis et al. (2023) to higher dimensions.
@@
Note added in v2: After the completion of this work, we became aware of Amalberti et al. (2024), which performs a multi- pole expansion of the long-wavelength effective action for radiative sources in higher dimensions. The authors find a new set of Weyl-type moments, which coincide with the stress moments of a stationary object defined in this work.
@@ Appendix B Hartle-Thorne metric in harmonic coordinates
Once the coordinate transformation is fixed, we can find the metric in harmonic gauge at the chosen perturbative order by means of (112). We notice that the coordinate transformation has some gauge redundancies since there are coefficients which are unfixed. At 1PM and 2PM the redundancy is parameterized by two coefficients, namely $\mathcal{C}^{(R)}_{1,2,0}$ and $\mathcal{C}^{(R)}_{1,2,2}$. At 3PM a new redundancy arises by means of the unfixed coefficients $\mathcal{C}^{(R)}_{3,0,0}$ and $\mathcal{C}^{(R)}_{3,2,2}$. While the full 3PM result is given in the attached Mathematica notebook Anc , at 1PM the explicit expression of the Hartle-Thorne metric in harmonic gauge reads
@@ Appendix C Myers-Perry black holes in $D=5$ in harmonic coordinates
As we did in the Hartle-Thorne case, replacing Eq. (123) in the harmonic equation, one is able to fix the coefficients of the coordinate transformation. In this case the equivalent of Eq. (115) for the Myers-Perry solution is way more involved, and since there is no physical information in it, we provide it in the attached Mathematica notebook Anc . However, it is worth mentioning that coherently to what we know from the amplitude calculation, the coordinate transformation develops two independent gauge redundancies at orders $O(Ga^{2})$, and two more at orders $O(G^{2}a^{0})$ and $O(G^{2}a^{2})$. In particular the redundancies at 2PM are related to the logarithmic terms and are the classical counterpart of the higher-loop vertex in Eq. (51).
Finally, we report the 3PM Myers-Perry metric in harmonic coordinates in the attached Mathematica file Anc , giving here for the sake of clarity only the expression at 1PM in the case in which $\mathcal{C}_{1,2,0}^{(R)}=0$ and $\mathcal{C}_{1,2,2}^{(R)}=-2/9$ such that $g_{R\Theta}=0$, resulting in the simplified expression
@@ Appendix D Black rings with one angular momentum in harmonic coordinates
Finally, considering the explicit expression of Eq. (130) (reported in the Mathematica file Anc ), the black ring metric in harmonic coordinates (in a suitable gauge), in which we are neglecting quadrupoles that are not spin-induced, reads